### v2

@@ V Conclusion
In this work, we introduced a multi-step approach to automated construction of a KG of nuclear fusion energy. By leveraging the inference power of LLMs, we extracted entities and relationships from 8358 scientific abstracts in the field creating a KG that effectively represents its domain knowledge. This comprehensive representation is not only beneficial for data storage but also serves as a foundation for the development of a KG-RAG machine. Our results highlight the system’s abstraction capabilities over a complex domain knowledge, demonstrating its potential for enhancing information retrieval and knowledge representation in specialized fields.
@@ VI Acknowledgments
This work was supported by the Hartree National Centre for Digital Innovation, a collaboration between STFC and IBM. We also thank the collaboration Fusion Computing Lab between the UKAEA and the STFC Hartree Centre for its partial support to A. Agnello during this work. The author used ChatGPT (OpenAI) to assist with improving the clarity and linguistic expression of the manuscript. All ideas, logical structure, methodology, and conclusions are the result of the author’s own work.
@@ Appendix A Appendix A: list of category types
List of category types along with their corresponding definitions. While these definitions were not used in our analysis and during NER, we include them for clarity.
Entity Type Description
Concept A general idea or theoretical model that is used to understand, describe, or predict phenomena related to fusion processes.
Nuclear Fusion Experimental Facility A specialized research site designed to study and advance nuclear fusion as a viable energy source.
Nuclear Fusion Technique A method or approach designed to achieve and sustain nuclear fusion reactions, for example: Inertial fusion, magnetic confinement.
Nuclear Fusion Device Type Categories of nuclear reactors used in nuclear fusion experiments, for example: tokamak, stellarator.
Nuclear Fusion System Component Key structural or functional parts of a fusion device critical for operation, for example: wall, target, divertor.
Nuclear Fusion System Configuration Physical and operational setups that include detailed descriptions of the devices, systems, and parameters.
Experimental Apparatus Specialized equipment or subsystems used to support fusion experiments, typically external to the main fusion device. For example: vacuum system, cryogenic system, electronic components.
Physical Process Processes governing the interactions and transformations of physical systems.
Physics Entity Quantifiable characteristics that describe the state or behavior of physical systems.
Field Configuration Specific arrangement and intensity of magnetic and/or electric fields used to confine, control, and stabilize the plasma in fusion reactors.
Particle Subatomic particles.
Chemical Element or Compound An element is a pure substance consisting of one type of atom, distinguished by its atomic number. A compound is made of different elements bonded together.
Plasma Property Defines specific attributes or states of plasma in fusion experiments, for example: plasma states and modes (H-mode, edge-localized modes (ELMs)), plasma temperature and density.
Plasma Event Distinct events that occur within plasma during nuclear fusion experiments, for example: disruption, thermal quenching.
Plasma Region Focuses on spatial divisions and areas within a plasma system, for example: core, edge, scrape-off layer.
Plasma Dynamics and Behavior Dynamic processes and interactions that occur within the plasma, for example: plasma transport, turbulent diffusion, magnetohydrodynamic (MHD).
Detection and Monitoring Systems Instruments and tools used to measure and analyze plasma behavior and reactor performance.
Control Systems Systems designed to actively regulate and maintain optimal conditions within the reactor.
Theory and Calculation Theoretical models, frameworks, or computational methods explaining or predicting fusion-related phenomena.
Software and Simulation Computational tools or software for simulating and analyzing fusion systems.
Time Reference A specific temporal reference that defines when an event occurred, such as a particular date, year.
Country and Location Geographical regions (such as countries or cities) or specific places and locations.
Facility or Institution Organizations, laboratories, or research establishments supporting fusion studies.
Person Individuals.
Safety Feature and Regulatory Standard Mechanisms, standards, guidelines ensuring safe operations in nuclear fusion environments.
Database Organized collection of structured information or data.
Scientific Publication and Citation Peer-reviewed written document presenting the results of scientific research or a reference to such a document used to acknowledge its contribution.
Research Field A specialized area of study.

### v3

@@ V Conclusion
In this work, we developed a modular, generalizable, and domain-agnostic multi-step approach to automated construction of a KG. To demonstrate feasibility of our approach, we applied it to the nuclear fusion domain, where minimal specific adaptations (like acronym expansion, and category selection) were required. In our pipeline, we leveraged the inference power of LLMs to extract entities and relationships from 8358 scientific abstracts in the field, creating a KG that effectively represents its domain knowledge. This comprehensive representation is not only beneficial for data storage but also serves as a foundation for the development of a KG-RAG system. Our results highlight the system’s abstraction capabilities over a complex domain knowledge, demonstrating its potential for enhancing information retrieval and knowledge representation in specialized fields.
@@ VI Acknowledgments
The authors used generative AI to improve the clarity of the manuscript. All ideas, logical structure, methodology, and conclusions are the result of the authors’ own work.
@@ Appendix A Appendix A: list of category types
List of the 28 category types used in our analysis and during NER: Concept, Nuclear Fusion Experimental Facility, Nuclear Fusion Technique, Nuclear Fusion Device Type, Nuclear Fusion System Component, Nuclear Fusion System Configuration, Experimental Apparatus, Physical Process, Physics Entity, Field Configuration, Particle and Subatomic particles, Chemical Element or Compound, Plasma Property, Plasma Event, Plasma Region, Plasma Dynamics and Behavior, Detection and Monitoring Systems, Control Systems, Theory and Calculation, Software and Simulation, Time Reference, Country and Location, Facility or Institution, Person, Safety Feature and Regulatory Standard, Database, Scientific Publication and Citation, Research Field.